OpenStudy (anonymous):

for what values of k is the locus of P(x,y) PF'-PF=k a hyperbola, pair of rays, null set.

OpenStudy (anonymous):

k must be >0

OpenStudy (anonymous):

for what

OpenStudy (anonymous):

I'm still having a mental block with this:- If P is on the hyperbola and the F's are the foci then k is 2a (ie constant) by definition

OpenStudy (anonymous):

what happens to the hyperbola when 2a=2c

OpenStudy (anonymous):

You hace c^2 = a^2-b^2 For c to equal a then b must be 0

OpenStudy (anonymous):

so the hyperbola becomes a line segment along x axis.. yeah?

OpenStudy (anonymous):

No, it is just not a hyperbola

OpenStudy (anonymous):

so do u get my queston now? :x

OpenStudy (anonymous):

No, a hyperbola (centred on (0,0) is x^2/a^2-y^2/b^2 = 1 (b cannot be 0)

OpenStudy (anonymous):

the locus doesnt have to be a hyperbola? for what values of k is the locus of P(x,y) PF'-PF=k a hyperbola, pair of rays, null set.

OpenStudy (anonymous):

a>b>0 is part of the definition of a hyperbola

OpenStudy (anonymous):

In your question, k=2a and is constant

OpenStudy (anonymous):

I think you are trying to ask a different question.

OpenStudy (anonymous):

but when k=2a=2c, hyperbola does not exist

OpenStudy (anonymous):

a=c is not defined for hyperbola

OpenStudy (anonymous):

k>c for locus to be hyperbola. what happens when k=c? should be a pair of rays

OpenStudy (anonymous):

Why are you making up definitions for hyperbola, there are standard definitions to use.

OpenStudy (anonymous):

Any other definition will not be a hyperbola

OpenStudy (anonymous):

im not.. it was a question my test which asked that. it deosnt have to be a hyperbola

OpenStudy (anonymous):

Well, that's a different question....... But you are using P(x,y) which is the usual notation for a point on a hyperbola and F, the usual notation for a focus.

OpenStudy (anonymous):

ie you are assuming a hyperbgola to start with.

OpenStudy (anonymous):

how about.. when k=2c for sum of distances, the locus is a line segment